# Speech Recognition on TV Series with Video-guided Post-Correction

### v1

@@ Abstract
Automatic Speech Recognition (ASR) has achieved remarkable success with deep learning, driving advancements in conversational artificial intelligence, media transcription, and assistive technologies. However, ASR systems still struggle in complex environments such as TV series, where overlapping speech, domain-specific terminology, and long-range contextual dependencies pose significant challenges to transcription accuracy. Existing multimodal approaches fail to correct ASR outputs with the rich temporal and contextual information available in video. To address this limitation, we propose a novel multimodal post-correction framework that refines ASR transcriptions by leveraging contextual cues extracted from video. Our framework consists of two stages: ASR Generation and Video-based Post-Correction, where the first stage produces the initial transcript and the second stage corrects errors using Video-based Contextual Information Extraction and Context-aware ASR Correction. We employ the Video-Large Multimodal Model (VLMM) to extract key contextual information using tailored prompts, which is then integrated with a Large Language Model (LLM) to refine the ASR output. We evaluate our method on a multimodal benchmark for TV series ASR and demonstrate its effectiveness in improving ASR performance by leveraging video-based context to enhance transcription accuracy in complex multimedia environments.
keywords:
speech recognition, post-correct, TV series, video-large multimodal model, large language model
footnotetext: *Equal Contribution.
@@ 1 Introduction
Automatic Speech Recognition (ASR) has achieved remarkable success with the advent of deep learning, driving advancements in applications such as conversational AI, media transcription, and assistive technologies [1, 2, 3]. Despite their success, ASR systems still struggle in complex real-world scenarios, particularly in noisy environments [4], domain-specific conversations [5], and multimedia content such as TV series or movies [6]. These scenarios often involve overlapping speech, domain-specific terminology, and contextual dependencies, making it difficult for ASR systems to rely solely on audio signals to achieve optimal performance.
To address the above challenges, leveraging multimodal contextual information has emerged as a promising direction [7]. However, the existing work focus on utilizing image information to correct the ASR output, but overlooks the information from the video. More importantly, ASR on TV series [8, 9, 10] plays a particularly crucial role, as it directly impacts the viewing experience by catering to audience needs, enhancing accessibility, and improving content engagement. Meanwhile, the rich context information from TV series videos can be utilized for the correct ASR output.
In particular, video serves as a powerful complementary modality, providing rich contextual cues that speech alone cannot convey. Scene elements—such as objects, gestures, and environmental settings—offer critical information for resolving ambiguous words and enhancing comprehension, particularly in noisy conditions. Additionally, character movements provide valuable speaker cues, helping to mitigate diarization errors. Furthermore, the long-range dependencies inherent in TV series and movies—where earlier scenes influence later dialogue—can significantly benefit from video-based context, aiding in maintaining coherence and accurately recognizing named entities.
@@
To fully utilize the rich context information from the TV Series video, we proposed a novel multimodal post-correction that refines ASR outputs by leveraging contextual content extracted from video. Our framework consists of two stages: 1) ASR Generation and 2) Video-guided Post-Correction (VPC). The first stage aims to generate ASR output with an ASR model. The later stage is for correcting errors in the generated transcript based on the video. Specifically, it consists of two main components: Video-based Contextual Information Extraction and Context-aware ASR Correction. Video-based Contextual Information Extraction utilizes the advanced Video-Large Multimodal Model (VLMM) to extract key context information with our devised prompts. Context-aware ASR Correction aims to correct errors in the generated transcript using the Large Language Model (LLM) and extracted content by VLMM. We evaluate our approach on a multimodal TV series dataset Violin [11] and ASR models wav2vec 2.0111https://huggingface.co/facebook/wav2vec2-large-960h., HuBERT [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. and WavLM [^3] https://huggingface.co/patrickvonplaten/wavlm-libri-clean-100h-large.. The experiments demonstrate that our method can significantly reduce the Word Error Rate (WER), such as lowering it from 45.5% to 35.4% for WavLM.
@@
Our contribution can be summarized as follows, 1) To the best of our knowledge, we are the first one to propose to post-correct for errors in ASR with additional information on video modality. 2) We propose a novel video-guided post-correction method, which can leverage advanced VLMM to extract visual information and LLM to correct the errors in ASR outputs. And 3) We conduct extensive experiments on the Violin dataset and ASR models. The experimental results show the effectiveness of our method.
@@ 2 Related Work
Since Language Models (LMs) [12] achieved success on various sequence modeling tasks, LMs play a fundamental role in ASR by capturing linguistic context to improve recognition accuracy [13, 14]. In traditional ASR frameworks, LMs assist in both decoding and rescoring by modeling the probability of word sequences [15, 16, 17]. Furthermore, the development of Pre-trained Language Models (PLMs) [18, 19] has significantly advanced language modeling in ASR, moving beyond traditional n-gram models [20] to deep contextual embeddings. Unlike early models [12, 17] focused solely on word-level probabilities, PLMs enable richer semantic representations and have achieved superior performance across various ASR benchmarks [21, 22]. While these improvements address many challenges in ASR decoding, recognition errors persist due to acoustic variability, domain mismatches, and noisy environments. This led to increased research interest in dedicated ASR correction models, evolving from statistical machine translation techniques to modern neural architectures designed to refine outputs after initial decoding. Early correction models employed Statistical Machine Translation (SMT) techniques trained on parallel corpora of ASR hypotheses and corrected transcripts [23]. The introduction of neural sequence-to-sequence models enabled more accurate corrections through large-scale synthetic corpora generated using text-to-speech (TTS) systems [24]. To address latency concerns, non-autoregressive transformer models were introduced, balancing speed and accuracy [25].
However, real-world ASR tasks often occur in multimodal contexts, where additional information from video and audio can enhance error correction. Therefore, we propose a novel multimodal post-correct framework that integrates visual context from video data to improve the quality of ASR outputs.
Figure 1: Overview of the proposed Video-Guided Post-Correction (VPC) method. The framework integrates video-based contextual information to refine ASR outputs.
@@ 3 Task Formulation
We first introduce the related dataset, then formulate the ASR on TV Series task, and finally introduce the post-correct task.
Dataset. Suppose we have a training dataset $\mathcal{D}$ composed of $N$ samples, i.e., $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{N}\}$. Each sample $d_{i}=\{A_{i},V_{i},Y_{i}\}$, where $A_{i}=\{a_{1}^{i},a_{2}^{i},\cdots,a_{N_{a_{i}}}^{i}\}$ denotes the input audio segmented into $N_{a_{i}}$ frames, $V_{i}$ represents the TV series video input, and $Y_{i}=\{y_{1}^{i},y_{2}^{i},\cdots,y_{N_{y_{i}}}^{i}\}$ denotes the target transcription text consisting of $N_{y_{i}}$ tokens.
ASR on TV Series. Based on these training samples, the target of ASR is to learn an ASR model $\mathcal{F}$ that maps the given audio input to its corresponding transcription as follows:
@@
where $\Theta$ represents the set of parameters to be learned for the ASR model $\mathcal{F}$, and $\hat{Y}_{i}$ is the predicted transcription of the $i$-th sample.
Video-guided ASR Post-correct. Based on the generated transcription from the ASR model, the video post-correct task aims to correct errors in the generated transcription as follows,
@@
where $\bar{Y}_{i}$ is the corrected transcription based on the $\hat{Y}_{i}$ and video $V_{i}$. $\mathcal{M}$ is our devised post-correct method.
@@ 4 Methodology
We propose a novel multimodal post-correct framework that integrates visual context from video data to improve the quality of ASR outputs. The framework consists of two main stages: ASR Generation Stage and Video-guided Post-Correction Stage.
@@ 4.1 ASR Generation
The large audio foundation models, wav2vec 2.0, HuBERT and WavLM, have achieved compelling success in several ASR datasets, such as Libri-Light [26], TIMIT [27], and Librispeech[28], however, they still struggle with named entity misrecognition errors, phonetic approximation errors, and homophones errors. For example, when transcribing TV series dialogues, it may misrecognize Joey Tribbiani as Joey Tribbyany. This error occurs because the model lacks sufficient world knowledge to correctly infer the proper spelling of named entities, especially those that are uncommon or do not frequently appear in the training data. Therefore, we devised a video-guided post-correction method to correct the output for ASR models. In this stage, we feed the ASR model with audio and get the corresponding generated transcript.
4.2 Video-guided Post-Correction
To enhance ASR output quality, we propose a Video-guided Post-Correction (VPC) method that leverages contextual information extracted from video content. The framework is illustrated in Figure 1, and consists of two main components: Video-based Contextual Information Extraction and Context-aware ASR Correction.
@@ 4.2.1 Video-based Contextual Information Extraction
Considering that the videos usually contain rich contextual information (e.g., scenes, actions, and characters in movies), we propose to extract visual information as additional information to correct the errors in the generated transcript. Since the VLMM [29, 30, 31] have demonstrated impressive performance in tasks such as video understanding, caption generation, and video-based question answering by integrating visual and textual data, we resort to the VLMM to capture the rich semantic context from the video.
Specifically, we extract the fine-grained information from the video with a Question-Answering (QA) format. We devised two questions to fully mine the information from the video: 1) TV Show Recognition and 2) Fine-grained Video Description. The potential reason we devise these two questions is: 1) TV show name can be utilized to retrieve knowledge related to the TV show, correcting the related error in the generated transcript, such as character name error; 2) ASR models sometimes generate errors which are unrelated to the video content or topic, and fine-grained video description could be used to correct these error information.
@@ 4.2.2 Context-aware ASR Correction
In this stage, we aim to correct the ASR outputs using the extracted visual context information $C_{1}$ and $C_{2}$. A straightforward approach would be to train a dedicated model for the correction task; however, this requires extensive data collection and model training, which is both time-consuming and resource-intensive. Recently, Large Language Models (LLMs) have shown excellent performance in language understanding and generation, therefore, we resort to LLMs to correct errors in transcripts with context information.
Specifically, we feed the LLM with a task instruction, the extracted context information $C_{1}$ and $C2$ from the input video as follows,
@@
To ensure the quality of the correction, we train $\mathcal{G}$ using a supervised loss that minimizes the word error rate (WER) between the corrected output $\hat{Y}$ and the ground truth transcription.
@@ 5.1 Experimental Setup
To ensure the focus on ASR, we create a Violin-TV subset by selecting clips from four TV series in which English is the primary spoken language: Friends, Desperate Housewives, How I Met Your Mother, and Modern Family. The subset comprises 10,003 TV clips, including the corresponding audio, transcript, and video. The average duration of the subset is 32.4 seconds, with an average speech density of 74.9%. With a total duration of 90.027 hours, we allocate a 72-hour training set with 7,983 clips, a 9-hour validation set with 1,007 clips and a 9-hour testing set with 1,013 clips. We use the original transcript in the dataset as the ground-truth transcript.
@@
Table 1: The performance (WER) comparison of models on the Violin-TV subset. w/ 1-hour FT Data and w/ 72-hour FT Data indicate the baseline models fine-tuned on 1-hour and 72-hour ASR data, respectively.
Model Setting w/ VPC WER
wav2vec 2.0 w/o FT Data 45.2
40.8
w/ 1-hour FT Data 33.3
30.3
w/ 72-hour FT Data 27.1
25.2
HuBERT w/o FT Data 40.0
35.2
w/ 1-hour FT Data 28.4
26.4
w/ 72-hour FT Data 24.4
23.3
WavLM w/o FT Data 45.5
35.4
w/ 1-hour FT Data 30.8
28.3
w/ 72-hour FT Data 24.4
22.3
We apply our VPC method to the generated transcripts from state-of-the-art models to assess its effectiveness in correcting transcriptions. In our VPC method, we apply VideoLLaMA2 [29] as our VLMM, and GPT-4o as our LLM. For the evaluation metric, we use WER.
We use pre-trained large speech foundation models wav2vec 2.0, HuBERT, and WavLM, as our baselines (around 317M parameters for each one). To verify the effectiveness of our method in extensive sceneries, we further fine-tune baselines with two different settings in terms of data size: 1) FT 1-hour: The model is fine-tuned on 1-hour subset randomly sampled from the training data of Violin-TV. 2) FT 72-hour: The model is fine-tuned on all the training data in Violin-TV. We further apply our method to model variants of the two settings.
@@ 5.2 Results Analysis
We present the WER results of different models on the Violin-TV subset in Table 1. From this Table, we have several observations: 1) A stronger original model tends to yield better performance with the same amount of fine-tuning data. For instance, after just one hour of fine-tuning, Wav2Vec 2.0 achieves a WER of 37.0%, while HuBERT and WavLM attain WERs of 33.0% and 41.8%, respectively, highlighting the impact of pre-trained model quality on adaptation efficiency. 2) For a given model, increasing the amount of fine-tuning data leads to improved performance. For example, fine-tuning with 72 hours of data reduces the WER of the three models to 29.1%, 26.3%, and 29.7%, respectively, demonstrating the positive correlation between data volume and model effectiveness. 3) Our method consistently improves the performance of different models. Regardless of the baseline architecture, applying our approach leads to lower WER across Wav2Vec 2.0, HuBERT, and WavLM, demonstrating its effectiveness in enhancing diverse ASR models. And 4) Our method also improves performance under different fine-tuning data settings for the same model. Whether using 1 hour or 72 hours of fine-tuning data, our approach consistently reduces WER, showing its robustness in leveraging varying amounts of labeled data to enhance ASR accuracy.
@@ 5.3 Case Study
We have shown a case about how our method corrects the character name error in wav2vec 2.0 w/ 72-hour FT Data in Figure 1. We further show another case study about comparison between results generated by WavLM w/ 72-hour FT Data and its corresponding variant augmented with our method. In this sample, GPT-4o revise the error words “a be hi hat” into the correct word “a beehive”. This is because the video content provides additional information to facilitate the LLM to revise the potential error by reasoning. This further verify the effectiveness of our method.
Figure 2: Comparison between results generated by WavLM w/ 72-hour FT Data and its corresponding variant augmented with our method. The red fonts are the highlighted error words. The green fonts are the highlighted useful information or corrected errors.
@@ 6 Conclusion
In this work, we introduced a novel video-guided post-correction (VPC) method to enhance ASR performance in complex multimedia environments, particularly in TV series. Unlike existing approaches, our method effectively leverages Video-Large Multimodal Models (VLMMs) to extract rich contextual information from videos and integrates it with Large Language Models (LLMs) to refine ASR outputs. Experimental results on a multimodal TV series benchmark demonstrate the effectiveness of our method in reducing ASR errors and enhancing overall transcription quality. Moving forward, our work paves the way for more robust multimodal ASR correction strategies that integrate deeper video understanding, making ASR systems more adaptable and reliable in real-world multimedia applications.

### v2

# SPEECH RECOGNITION ON TV SERIES WITH VIDEO-GUIDED POST-ASR CORRECTION

@@ Abstract
Automatic Speech Recognition (ASR) has achieved remarkable success with deep learning, driving advancements in conversational artificial intelligence, media transcription, and assistive technologies. However, ASR systems still struggle in complex environments such as TV series, where multiple speakers, overlapping speech, domain-specific terminology, and long-range contextual dependencies pose significant challenges to transcription accuracy. Existing approaches fail to explicitly leverage the rich temporal and contextual information available in the video. To address this limitation, we propose a Video-Guided Post-ASR Correction (VPC) framework that uses a Video-Large Multimodal Model (VLMM) to capture video context and refine ASR outputs. Evaluations on a TV-series benchmark show that our method consistently improves transcription accuracy in complex multimedia environments.
Index Terms—  post-ASR correction, video-audio, video-large multimodal model
@@ 1 Introduction
Automatic Speech Recognition (ASR) has achieved remarkable success with the advent of deep learning, driving advancements in applications such as conversational AI, media transcription, and assistive technologies [1]. Despite their success, ASR systems still struggle in complex real-world scenarios, particularly in noisy environments [2], domain-specific conversations [3], and multimedia content such as TV series or movies [4]. These scenarios often involve overlapping speech, domain-specific terminology, and contextual dependencies, making it difficult for ASR systems to rely solely on audio signals to achieve optimal performance.
To address the above challenges, leveraging multimodal contextual information has emerged as a promising direction [5]. However, the existing works focusing on utilizing image information to correct the ASR output [6] lack full exploitation of video information [7]. ASR on TV series [8, 9] is particularly important for accessibility and engagement. In particular, video serves as a powerful complementary modality, providing rich contextual cues that speech alone cannot convey. For example, the video-based context, such as scene elements, character movements, and long-range dependencies inherent in TV series and movies, can significantly aid in maintaining coherence and provide more useful information for recognizing named entities.
@@
To fully utilize the rich context information from the TV series video, we propose a novel training-free multimodal post-ASR correction (VPC) framework that refines ASR outputs by leveraging contextual content extracted from videos. Our framework consists of two stages: 1) ASR Generation, where an ASR model will transcribe the audio; 2) Video-guided Post-ASR Correction. The latter stage is for correcting errors in the generated transcript based on the video. Specifically, it consists of two main components: I. The Video-based Contextual Information Extraction module, utilizing the advanced Video-Large Multimodal Model (VLMM) to extract key context information with our devised prompts; II. The Context-aware ASR Correction module aims to correct errors in the generated transcript using the Large Language Model (LLM) and extracted visual context.
We evaluate our approach on a multimodal TV series dataset Violin [10] with state-of-the-art self-supervised speech models (wav2vec 2.0 [11], HuBERT [12], WavLM [13]), and Conformer [14]. The experiments demonstrate that our method can significantly reduce the Word Error Rate (WER) by providing necessary visual context, yielding a relative 20.75% improvement on the Violin dataset [10] for WavLM.
@@
Our contribution can be summarized as follows, 1) To the best of our knowledge, we are the first ones to propose to post-correct for errors in ASR with additional information on video modality. 2) We propose a novel video-guided post-ASR correction method, which can leverage advanced VLMM to extract visual information and LLM to correct the errors in ASR outputs. And 3) We conduct extensive experiments on the Violin dataset and ASR models. The experimental results show that our proposed method provides meaningful visual context that can help improve ASR accuracy under complex multimodal scenarios.
@@ 2 Related Work
Recent post-ASR correction [15] aims at correcting the generative errors of a pretrained ASR model using LLMs, with training-free strategy [16] or light additional training [17]. However, real-world ASR tasks often occur in multimodal contexts, where additional information from video can enhance error correction. Audio-visual models [6, 7, 18] improve recognition by jointly encoding visual and audio features. In contrast, the role of video context in post-ASR correction remains underexplored, despite its potential to mitigate errors that persist after decoding. While these models have demonstrated strong performance, an important research gap remains: the potential role of visual information in further improving ASR performance. VLMMs [19, 20, 21] have excelled in tasks such as video understanding, captioning, and video-based question answering, but their application to ASR correction remains unexplored. Therefore, different from the existing work, we propose a novel training-free multimodal post-ASR correct framework, which integrates visual context from video data to improve the quality of ASR outputs.
Fig. 1: Overview of the proposed Video-Guided Post-ASR Correction (VPC) method. The framework integrates video-based contextual information to refine ASR outputs.
@@ 3 Task Formulation
We first introduce the related dataset, then formulate the ASR on TV Series task, and finally introduce the post-ASR correction task. 1) Dataset. Suppose we have a training dataset $\mathcal{D}$ composed of $N$ samples, i.e., $\mathcal{D}=\{d_{1},d_{2},\cdots,d_{N}\}$. Each sample $d_{i}=\{A_{i},V_{i},Y_{i}\}$, where $A_{i}$ denotes the input audio segmented into $N_{a_{i}}$ frames, $V_{i}$ represents the TV series video input, and $Y_{i}$ denotes the target transcription text consisting of $N_{y_{i}}$ tokens. 2) ASR on TV Series. Based on these training samples, the target of ASR is to learn an ASR model $\mathcal{F}$ that maps the given audio input to its corresponding transcription as follows:
@@
where $\Theta$ represents the set of parameters to be learned for the ASR model $\mathcal{F}$, and $\hat{Y}_{i}$ is the predicted transcription of the $i$-th sample. 3) Video-guided post-ASR correction. Based on the generated transcription from the ASR model, the video-based correction task aims to identify and repair errors in the generated transcription as follows,
@@
where $\bar{Y}_{i}$ is the corrected transcription based on the $\hat{Y}_{i}$ and video $V_{i}$. $\mathcal{M}$ is our devised post-ASR correction method.
@@ 4 Methodology
We propose a novel multimodal post-ASR correction framework that integrates visual context from video data to improve the quality of ASR outputs. The framework consists of two main stages: ASR Generation Stage and Video-guided Post-ASR Correction Stage.
@@ 4.1 ASR Generation
Despite the compelling success recent ASR systems have achieved in several datasets [22, 23, 24], they still struggle with named entity misrecognition errors, phonetic approximation errors, and homophone errors. For example, when transcribing TV series dialogues, it may misrecognize Joey Tribbiani as Joey Tribbyany. This error occurs because the model lacks sufficient world knowledge to correctly infer the proper spelling of named entities, especially those that are uncommon or do not frequently appear in the training data. Therefore, we devised a video-guided post-ASR correction method to correct the output for ASR models. In this stage, we feed the ASR model with audio and get the corresponding generated transcript $\hat{Y}$.
4.2 Video-guided Post-ASR Correction
To enhance ASR output quality, we propose a Video-guided Post-ASR Correction (VPC) method that leverages contextual information extracted from video content. The framework is illustrated in Figure 1, and consists of two main components: Video-based Contextual Information Extraction and Context-aware ASR Correction.
@@ 4.2.1 Video-based Contextual Information Extraction
Given VLMMs’ impressive performance in video understanding, we resort to VLMM to capture the rich semantic context (e.g., scenes, actions, and characters) in video. Specifically, we extract the fine-grained information from the video with a Question-Answering (QA) format. We devised two questions to fully mine the information from the video: 1) TV Show Recognition and 2) Fine-grained Video Description. The potential reason we devised these two questions is: 1) the TV show name can be utilized to retrieve knowledge related to the TV show, such as character names, and the QA-based information extraction method could mine this kind of information from videos; 2) ASR models sometimes generate errors that are unrelated to the video content or topic, and fine-grained video description could be used to correct such errors.
@@ 4.2.2 Context-aware ASR Correction
In this stage, we aim to correct the ASR outputs using the extracted visual context information $C_{1}$ and $C_{2}$. A straightforward approach would be to train a dedicated model for the correction task; however, this requires extensive data collection and model training, which are both time-consuming and resource-intensive. Recently, Large Language Models (LLMs) have shown excellent performance in language understanding and generation; we resort to LLMs to correct errors in transcripts with context information.
Specifically, we feed the LLM with a task instruction, the extracted context information $C_{1}$ and $C_{2}$ from the input video as follows,
@@
Model Raw ASR Post-ASR Correction
GPT (w/o visual context) Raw-Impro↑ VPC Raw-Impro↑ GPT-Impro↑
wav2vec2-large 29.17 29.28 -0.38% 25.36 13.06% 13.39%
HuBERT-large 26.40 25.73 2.54% 23.27 11.86% 9.56%
WavLM-large 29.83 29.45 1.27% 23.64 20.75% 19.73%
Conformer-large 22.66 22.14 2.29% 20.97 7.46% 5.28%
Table 1: WER comparison of models on the Violin-TV subset. GPT uses no visual context prompt, while VPC augments GPT with VLMM-based visual context. Raw-Impro and GPT-Impro are the relative WER improvement between our VPC and Raw ASR and GPT (w/o visual context), respectively.
@@ 5.1 Experimental Setup
To ensure the focus on ASR, we create a Violin-TV subset by selecting clips from only TV series in which English is the primary spoken language. The subset comprises 10,003 TV clips, including the corresponding audio, transcript, and video. The average duration of the subset is 32.4 seconds, with an average speech density of 74.9%. With a total duration of 90.027 hours, we allocate a 72-hour training set with 7,983 clips, a 9-hour validation set with 1,007 clips and a 9-hour testing set with 1,013 clips. We use the original transcript in the dataset as the ground-truth transcript.
@@
To assess our VPC’s effectiveness in correcting transcription, we selected wav2vec 2.0, HuBERT and WavLM, and Conformer as our ASR models. All models are pretrained on Librispeech-960h [24]. We finetuned wav2vec 2.0, HuBERT and WavLM with CTC loss, and finetuned Conformer with RNN-T loss on the 72-hour training set. In our VPC method, we apply VideoLLaMA2 [19] as our VLMM, and GPT-4o as our LLM. For the evaluation metric, we use WER.
@@ 5.2 Results Analysis
We present the WER results of different models on the Violin-TV subset in Table 1. From this Table, we have several observations: 1) Proposed VPC constantly improves all models. With additional visual context, the proposed VPC achieves 20.75% relative WER reduction on WavLM-Large, 13.06% on wav2vec 2.0, 11.86% on HuBERT and 7.64% on Conformer-Large with RNN-T loss. This highlights that leveraging video information using QA and helping refine ASR output is effective and robust. 2) VPC is Training-free and generalizable. Our proposed VPC needs no further training with ASR models. The improvements across different models suggest that it is easy to integrate and generalize. 3) GPT-4o alone without visual information performs worse than our VPC method. Using GPT-4o directly yields negligible or even negative effects (e.g., –0.38% relative on wav2vec2-large) compared to the original ASR output. This indicates that text-only LLMs cannot reliably handle ASR correction in complex multimodal settings, where cross-modal dependencies and long-range context play an essential role.
Fig. 2: Comparison between results generated by WavLM and its corresponding variant augmented with VPC.
@@ 5.3 Case Study
We further show a case study about the comparison between results generated by WavLM and its corresponding variant augmented with our method. In this sample, our proposed model revises the error words “a be hi hat” into the correct word “a beehive”. This is because the video content provides additional information to facilitate the LLM to revise the potential error by reasoning. This further verifies the effectiveness of our method.
@@ 6 Conclusion
In this work, we introduced a novel video-guided post-ASR correction (VPC) method to enhance ASR performance in complex multimedia environments. By effectively leveraging Video-Large Multimodal Model (VLMM) to extract rich contextual information from videos and integrating it with Large Language Model (LLM), our method significantly reduces ASR errors on a multimodal TV benchmark. Experimental results demonstrate the effectiveness and generalizability of our method. Moving forward, our work paves the way for more robust multimodal ASR correction strategies that integrate deeper video understanding, making ASR systems more adaptable and reliable in real-world multimedia applications.